LBT SW Testing and Range Balancing:
UT 2011 Jan 28

Observer: DThompson Olga Kuhn and Andrew Rakich (Tucson)
Telescope Operator: DHuerta
Instrument Support: JMorris
Telescope Support: JLittle
SW support: CBiddick and MDeLaPena in Tucson

Summary

  • Problems running up LBC because of IP address change
  • Binocular guiding tests with LBC and BP8
  • Checked LUCI (range-balanced) collimation and pointing models.
  • Tested new WFS algorithm (DMiller)
  • Problems running IRTC through IRS, because of IP address change?

Details

Swapped to binocular LBC code (lbckill, change pointer, lbcstart). UI refused to come up (no permission). JHill suggested rebooting lbccontrol. Still refuses to start UI. Swapped back and it still has the same problem. You see the initial page, but when you try to start the UI you get:
Forbidden
You don't have permission to access /ui/index.html on this server.

and this did not work as LBTO or as dthompson.

This appears to be caused by the IP address/VLAN/subnet change made to the obsN workstations yesterday or today as the system appears to run fine from my laptop . We had to add the line "Allow from 192.168.58" to the /etc/httpd/httpd.conf file in the section controlling directory access to /home/lbccontrol/www/ui and then restart the web server to get the connection to work.

02:25 warmup LBC for 15 minutes...
02:40 turn on LBC...everything came up OK!
02:54 We had a "restart after crash" message but only the blue side came up. Did a "turn off" then "warmup" while Chris fixes some other problems coming from the IP address change.
03:02 "turn on" LBC...everything came up OK again!
03:03 Connect to LBT with right arm only...it worked!
03:06 dofpia, /redonly, /first seems to be doing something reasonable, but...
Processing: /newdata/lbcr.20110128.030806.fits
% Compiled module: GETIMAGE.
% Compiled module: FITS_OPEN.
% FITS_OPEN:  ERROR: Invalid header, no SIMPLE keyword
The image comes across with the detector data but no header at all. So iraf/idl tasks will not work. Trying additional tests with just the default collimation model (actually looks good: 0.8 arcsec, only a small amount of coma visible).

red test: 033001 - 034016 monocular guiding mount
blue test 034633 - 040032 monocular guiding mount
Fourth image had this error:
2011/01/28   04:00:10.566000   CAMERA    CCDCTRL    too few downloaded pixels (42466720 instead of 42467328) [ccdctrl.c:2696]  
but the exposure was done over automatically.

red+blue test binocular, guiding optics
red: 041223 -
blue: 041240 -
Guide on blue says it failed about halfway through the first exposure, failed again on the other B exposures. Red seems pretty stable.

05:20 Switched to AO5 to make sure things still work. TCS connection error. Chris made a change on the TCS side and we still had the same error, we did an lbckill/lbcstart and then a full startup from the UI. On startup the red camera gave:
2011/01/28   05:26:27.390687   CAMERA       RPC Client/Host communication failed retval:5 [src/camera/camera.c:549] 

I shut down the "other systems" and housekeeping, then did an lbckill/lbcstart and powered things back up. There was a problem on blue with it being "hinibit"ed, then when I tried turning them back on I got:
2011/01/28   05:34:32.045675   ROTATOR       GOYA no prompt return [src/rotator/rotator.c:2078]
58   2011/01/28   05:34:30.765626   ROTATOR       GOYA no prompt return [src/rotator/rotator.c:2078]
57   2011/01/28   05:34:29.021000   CAMERA    CCDCTRL    safetyoff completed
56   2011/01/28   05:34:29.021000   CAMERA    CCDCTRL    safetyoff has been forced
55   2011/01/28   05:34:28.990000   CAMERA    CCDCTRL    "FPGA not loaded" (code:48) controller error [ccdctrl.c:2469]
54   2011/01/28   05:34:27.881000   CAMERA    CCDCTRL    safetyoff has been forced
53   2011/01/28   05:34:27.865000   CAMERA    CCDCTRL    "CCD Controller fiber link failure" (code:38) controller error [ccdctrl.c:2463] 
But in fact all systems came back up. Image headers are populated and we can send presets, so it looks like AO5 still works. TCS Testing declared complete, turning it over to Andrew for further tests.

05:46 Run dofpia on both sides...blue failed to find pupils the first time, although it should not have been a problem. We sent -1000nm Z7 and tried again. RED converged in 1 iteration! Blue is heading in the right direction...converged! Running dofpia again while I look into a newline problem in lbcrangebal.cl in the BPbuild declaration.

PARK LUCI at 341 degrees with NO MASK in the fpu for mask exchange tomorrow !!

Test of LBC Range Balancing

* lbcrangebal in LBTtools.Observe package

Initial Test on 3 pointing stars at a different elevations:

75 deg elevation:

R 70549 B 70527 * Pointing Model Terms * IA -305.29050 IE -70.24410 NPAE 96.43150 CA -204.50290 AW 1.21960 AN 21.32600 TF 4.51840 TX 12.82410

lbcrangebal...

confirmation: R 70934 B 70908 * Pointing Model Terms * IA -305.29050 IE -34.03257 NPAE 96.43150 CA -166.07410 AW 1.21960 AN 21.32600 TF 4.51840 TX 12.82410

and mirror position:

R 70934:
MIRRORX =               -1.039 / Primary mirror X position [mm]
MIRRORY =               -0.841 / Primary mirror Y position [mm]
MIRRORZ =                0.035 / Primary mirror Z position [mm]
MIRRORRX=              -18.393 / Primary mirror RX rotation [arcsec]
MIRRORRY=               35.225 / Primary mirror RY rotation [arcsec]
MIRRORRZ=               -0.019 / Primary mirror RZ rotation [arcsec]

B 70908:
MIRRORX =                0.517 / Primary mirror X position [mm]
MIRRORY =                0.749 / Primary mirror Y position [mm]
MIRRORZ =               -0.120 / Primary mirror Z position [mm]
MIRRORRX=               23.688 / Primary mirror RX rotation [arcsec]
MIRRORRY=              -65.927 / Primary mirror RY rotation [arcsec]
MIRRORRZ=               -0.015 / Primary mirror RZ rotation [arcsec]

== 50 deg elevation: ACT0182 B 071540 (1038.88, 2909.9) [1040,2921] R 071605 (1091.6, 2908.1) [1080,2912]

lbcrangebal...

B 072649 R no red image???

B 72815 (1043.8,2933) R 72849 (1078,2920) lbcrangebal...

B 73226 (1040.4,2924) R 73256 (1078,2913)

== 25 deg elevation: ACT0097 B 73559; R 73628

lbcrangebal...
          Old      Delta      New
       -------   -------   -------
   CE  -35.016    -3.801   -38.817
   CA -165.276     3.873  -161.403

B 73840 (1039.8,2922.4) R 73910 (1080,2912)

R 73910:
MIRRORX =               -1.843 / Primary mirror X position [mm]
MIRRORY =               -1.132 / Primary mirror Y position [mm]
MIRRORZ =                1.846 / Primary mirror Z position [mm]
MIRRORRX=              -19.434 / Primary mirror RX rotation [arcsec]
MIRRORRY=               13.616 / Primary mirror RY rotation [arcsec]
MIRRORRZ=                0.021 / Primary mirror RZ rotation [arcsec]

B 73840:
MIRRORX =                1.158 / Primary mirror X position [mm]
MIRRORY =                0.944 / Primary mirror Y position [mm]
MIRRORZ =                1.828 / Primary mirror Z position [mm]
MIRRORRX=               27.998 / Primary mirror RX rotation [arcsec]
MIRRORRY=              -44.965 / Primary mirror RY rotation [arcsec]
MIRRORRZ=               -0.001 / Primary mirror RZ rotation [arcsec]

lbcrangebal script works

What is moving to cause Y/RX jumps? See IT #879

* Elevation 57: ACT0201: Collimated and then took images. Ran lbcrangebal and took confirmatory pair of images: B=75741 (1042.5,2921.8); R=75814 (1081.8,2915.3)

* Now move to zenith. Click collimate on both SX and DX. Then move to horizon.

* Now back to ACT0201: Collimate... did not see any significant change in mirror position.

* Moved back to zenith and then to horizon and back to ACT0201. Pair of images allowed. Collimating... again no significant change in coma. Failed to replicate the Y/RX jump. It would be worthwhile during the SciOps nights to make a range-bal copointed collimation mode:
  • slew to a pointing star,
  • run rangebal,
  • record Instrument Offsets.

Collecting data for LBC range-bal co-pointed collimation model

* ACT0305; elev=84deg*: B=83028 R=83105; lbcrangebal
these are deltas...
SX: dX = 0.002 dY = 0.093 dRX = 1.994 dRY = -0.037    
DX: dX = 0.044 dY = -0.059 dRX = -1.265 dRY = -0.939
 
SX: 1.357 1.249 26.8 -29.1
DX: -1.263 -1.128 -27.205 27.09

          Old      Delta      New
       -------   -------   -------
   CE  -38.687     6.525   -32.162
   CA -167.125     3.894  -163.231

* ACT0256; elev=71.9deg*: B=83535, R=83612; lbcrangebal
SX: dX = 0.055 dY = -0.051 dRX = -1.098 dRY = -1.171
DX: dX = -0.081 dY = 0.047 dRX = 1.011 dRY = 1.740

          Old      Delta      New
       -------   -------   -------
   CE  -32.162     0.765   -31.397
   CA -163.231     0.603  -162.628

* ACT0220; elev=60deg*: B=84055 R=84133; lbcrangebal
SX: dX = 0.023 dY = -0.032 dRX = -0.682 dRY = -0.503
DX: dX = -0.028 dY = 0.042 dRX = 0.894 dRY = 0.606
          Old      Delta      New
       -------   -------   -------
   CE  -31.397    -5.119   -36.516
   CA -162.628    -3.903  -166.531

* ACT0167; elev=46.6deg*: B=84444, R=84523; lbcrangebal
SX: dX = -0.026 dY = 0.047 dRX = 1.015 dRY = 0.568
DX: dX = 0.050 dY = -0.025 dRX = -0.542 dRY = -1.081
          Old      Delta      New
       -------   -------   -------
   CE  -36.516    -0.386   -36.902
   CA -166.531     1.481  -165.050

* ACT0150; elev=40deg*: B=84826, R=84906; lbcrangebal
SX: dX = 0.015 dY = 0.044 dRX = 0.950 dRY = -0.322
DX: dX = 0.003 dY = -0.028 dRX = -0.602 dRY = -0.066
          Old      Delta      New
       -------   -------   -------
   CE  -36.902     0.663   -36.239
   CA -165.050     0.639  -164.411

* ACT0116; elev=31.4deg*:: B=85148, R=85232; lbcrangebal
SX: dX = 0.109 dY = 0.087 dRX = 1.862 dRY = -2.333
DX: dX = -0.078 dY = -0.052 dRX = -1.119 dRY = 1.674
          Old      Delta      New
       -------   -------   -------
   CE  -36.239    -6.105   -42.344
   CA -164.411    -0.388  -164.799

* ACT0084; elev=20.8deg*: B=85539, R=85620; lbcrangebal
SX: dX = 0.399 dY = 0.158 dRX = 3.395 dRY = -8.562
DX: dX = -0.355 dY = -0.082 dRX = -1.765 dRY = 7.615
          Old      Delta      New
       -------   -------   -------
   CE  -42.344     1.463   -40.881
   CA -164.799     5.169  -159.629

* ACT0099; elev=25.9deg*: B=85930, R=90012; lbcrangebal
SX: dX = -0.226 dY = -0.105 dRX = -2.245 dRY = 4.842
DX: dX = 0.205 dY = 0.063 dRX = 1.345 dRY = -4.401
          Old      Delta      New
       -------   -------   -------
   CE  -40.881    -2.000   -42.881
   CA -159.629    -2.507  -162.136

* ACT0116; elev=31.6deg*: B=90234, R=90316; lbcrangebal
SX: dX = -0.182 dY = -0.076 dRX = -1.637 dRY = 3.901
DX: dX = 0.159 dY = 0.042 dRX = 0.898 dRY = -3.409
          Old      Delta      New
       -------   -------   -------
   CE  -42.881     2.271   -40.610
   CA -162.136     0.345  -161.791

* ACT0150; elev=40.9deg*: B=90544, R=90626; lbcrangebal
SX: dX = -0.132 dY = -0.122 dRX = -2.610 dRY = 2.839
DX: dX = 0.100 dY = 0.086 dRX = 1.835 dRY = -2.152
          Old      Delta      New
       -------   -------   -------
   CE  -40.610     1.954   -38.655
   CA -161.791    -1.002  -162.793

* ACT0167; elev=47.3deg*: B=90840, R=90920; lbcrangebal
SX: dX = 0.046 dY = -0.050 dRX = -1.073 dRY = -0.992
DX: dX = -0.064 dY = 0.034 dRX = 0.738 dRY = 1.378
          Old      Delta      New
       -------   -------   -------
   CE  -38.655    -0.701   -39.357
   CA -162.793    -3.443  -166.236

* ACT0203; elev=54deg*: B=91224, R=91219; lbcrangebal
SX: dX = 0.029 dY = -0.043 dRX = -0.928 dRY = -0.631
DX: dX = -0.046 dY = 0.028 dRX = 0.595 dRY = 0.989
          Old      Delta      New
       -------   -------   -------
   CE  -39.357     0.359   -38.998 
   CA -166.236    -0.429  -166.665

* ACT0239; 65.5deg*: B=91526, R=91522;
  • dofpia - only 2 pupils: B=92110, R=92106; lbcrangebal
SX: dX = 0.160 dY = -0.003 dRX = -0.069 dRY = -3.437
DX: dX = -0.125 dY = -0.027 dRX = -0.587 dRY = 2.692
          Old      Delta      New
       -------   -------   -------
   CE  -38.998     2.372   -36.625
   CA -166.665     5.120  -161.545

* ACT0256; elev=72deg*: B=92447, R=92443; lbcrangebal (09:26 UT network connection to obs4 hung...Dave running lbcrangebal)


10:00 Reconfigure for LUCI. Files luci.20110128.NNNN.fits
  • Restarted Telescope and GEIRS services before logging in as observer (Build switch to BP7)
  • All OK on /home/engineer/bin/showSupervisorStatus.sh
  • 0001-0002 2s darks in O2DCR mode, readnoise=11.51e-
  • 0003-0004 10s darks in MER mode, readnoise=5.49e-
  • Time is the same as on obs2

* Check existing collimation and pointing models*
  • Conditions are stable (SXM1 -1.8 Outside -1.9, stable over first half of night).
  • Collimated at EL~59, checked at EL~58, 31, 82 and pointing looks decently good!

LUCI Pupil image #0011, looks the same as before.

ARakich/DMiller working on new WFS algorithm

Using DMiller's IDL code to collimate:

10:53UT wfsc 01-10 BS9148, guide star 0. Work needs to be done on scale factors - focus, and other zernikes are oscillating.

11:35 moving to IRTC tasks.

Talking to camera returns "IRS is not running". Call SHooper to fix.

DX M2 does not appear to be "flattening" as expected. Runa is looking into it.

-- DavidThompson - 28 Jan 2011
Topic revision: r9 - 28 Jan 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback